a shunt motor is running off a 220 - V supply taking an armature current of 15A , the resistance of the armature circuit being 0.8ohms . calculate the value of the generated e.m.f . jf the flux were suddenly reduced by 10 percent , to what value would the armature current increase momentarily ?


1
Expert's answer
2021-03-10T05:53:13-0500

Given quantities:

"U = 220V" "I = 15A" "R = 0.8\varOmega"

The electromotive force (e.m.f) is defined as the energy per unit charge that converted from other forms of energy to electrical energy to move a charge across the whole circuit.

"E = \large\frac{W}{Q} = \frac{\frac{U^2t}{R}}{It}=\frac{U^2}{IR} = \frac{220*220}{15*0.8} =" 4033N

"\varPhi_2 = 0.9\varPhi_1" "X_L = wL"

"\large\frac{L_1I_1}{2} =0.9* \frac{L_2I_2}{2}"


"\frac{R_1I_1}{2w} =0.9* \frac{R_2I_2}{2w}"

"U = 0.9*0.8*I_2 \to I_2 = 306A"